OpenStudy (anonymous):

Barkley has a box shaped like a cylinder that he will use to package a ball for a gift. He wants to fill the remainder of the box with packing foam. Which formulas will he need to use to find the volume of packing foam he will need? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.)

OpenStudy (anonymous):

A. V=pir^2h B. V=1/3pir^2 C. V=4/3pir^3 D. V=1/3lwh

OpenStudy (anonymous):

@e.mccormick @amistre64 @bohotness @EclipsedStar